# True Zero-Shot Inference of Dynamical Systems Preserving Long-Term Statistics

### v1

@@ Abstract
Complex, temporally evolving phenomena, from climate to brain activity, are governed by dynamical systems (DS). DS reconstruction (DSR) seeks to infer generative surrogate models of these from observed data, reproducing their long-term behavior. Existing DSR approaches require purpose-training for any new system observed, lacking the zero-shot and in-context inference capabilities known from LLMs. Here we introduce DynaMix, a novel multivariate ALRNN-based mixture-of-experts architecture pre-trained for DSR, the first DSR model able to generalize zero-shot to out-of-domain DS. Just from a provided context signal, without any re-training, DynaMix faithfully forecasts the long-term evolution of novel DS where existing time series (TS) foundation models, like Chronos, fail – at a fraction of the number of parameters and orders of magnitude faster inference times. DynaMix outperforms TS foundation models in terms of long-term statistics, and often also short-term forecasts, even on real-world time series, like traffic or weather data, typically used for training and evaluating TS models, but not at all part of DynaMix’ training corpus. We illustrate some of the failure modes of TS models for DSR problems, and conclude that models built on DS principles may bear a huge potential also for advancing the TS prediction field.
@@ 1 Introduction
  • Accurate zero-shot DSR: DynaMix achieves strong generalization across diverse DS, eliminating the need for extensive fine-tuning while maintaining accuracy in reproducing attractor geometry and long-term statistics.
  • Multivariate information transfer: Due to its multivariate architecture, the model efficiently captures dependencies among multiple system dimensions, enabling accurate reconstruction of their coupled dynamics. Further, it is neither bound to a specific dimensionality, but can flexibly adapt to other dimensions through specific embeddings, nor to a specific context length.
  • Computational and parameter efficiency: The model achieves high performance with a very lightweight architecture, enabling orders of magnitude faster inference than other foundation models.
@@ 2 Related work
Large language models (LLMs) exhibit an impressive ability to infer patterns from prompts (context) and generalize to novel situations without retraining [10, 27, 21, 18], although the underlying reasons for this are still a matter of debate [49, 85, 90]. This inspired the development of general-purpose time series (TS) foundation models which could accurately forecast time series from a short segment provided ‘in-context’ [2, 26], without the need of task-specific fine-tuning [20]. One idea is to simply use pretrained LLMs directly as TS forecasters [34, 67, 75, 88, 79], but the inherent differences between textual and many temporal data, such as continuity, present significant challenges [78]. To overcome these, transformer-based architectures, similar in design to LLMs, such as Chronos [2] or Moirai [51], were specifically pretrained on a large corpus of time series data. Promising zero-shot forecasting capabilities have also been achieved with alternative architectural designs, such as Tiny Time Mixers [26] or state-space models like Mamba [5].
@@ 3.1 Model architecture
To enable zero-shot reconstruction of novel DS, here we develop a specific mixture-of-experts (MoE) architecture that can be pretrained across many diverse DS (Fig. 9), with different experts possibly specializing in different dynamical regimes. As a base model for the experts, we leverage a recent parameter-friendly RNN designed to yield topologically parsimonious and interpretable representations of DS it is being trained on, the Almost-Linear RNN (AL-RNN; [7]):
@@
The selection of AL-RNN experts is achieved through a gating network (Fig. 2), which receives as inputs the (generally multivariate) context time series $\bm{C}=\{\bm{c}_{t}\}\in\mathbb{R}^{N\times T}$ as well as the current latent state $\bm{z}_{t}$. Both are passed into a state attention mechanism defined as
$\bm{w}^{att}_{t}=\sigma\left(\frac{\left|\bm{C}-\left(\bm{Dz}_{t}+\bm{\epsilon}\right)\bm{1}_{T}^{\top}\right|^{\top}\bm{1}_{N}}{T_{\text{att}}}\right)\in\mathbb{R}^{T}\;,$ (3)
where $\bm{D}\in\mathbb{R}^{N\times M}$ is a learnable matrix which maps the latent state into observation space, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{\Sigma})$ exploration noise with a learnable covariance $\bm{\Sigma}$, $T_{\text{att}}$ a learnable temperature parameter, $\bm{1}_{\{T,N\}}$ column vectors of ones of length $T,N$, respectively, and $\sigma(\cdot)$ the softmax returning normalized weights. Hence, this mechanism computes attention weights based on some distance between projected latent states and actual context observations. At the same time, the context signal $\bm{C}$ is processed by a CNN, yielding temporal features $\tilde{\bm{C}}\in\mathbb{R}^{N\times T}$. These features are then weighted with the attention weights $\bm{w}^{att}_{t}$, and together with the current state $\bm{z}_{t}$ further processed by a multi-layer perceptron (MLP) followed by a softmax, yielding a set of expert weights at any time $t$:
$\bm{w}^{exp}_{t}=\sigma\left(\frac{\text{MLP}(\tilde{\bm{C}}{\bm{w}^{att}_{t}},\bm{z}_{t})}{T_{\text{exp}}}\right)\in\mathbb{R}^{J}\;,$ (4)
where $T_{\text{exp}}$ is another learnable temperature parameter. The forward-iterated latent states $\bm{z}_{t+1}^{j}$ of the individual experts $j\in\{1,...,J\}$ are then weighted by $\bm{w}^{exp}_{t}$ to yield the next time step prediction $\bm{z}_{t+1}=\sum_{j=1}^{J}w^{exp}_{j,t}\cdot\bm{z}_{t+1}^{j}$. Fig. 2 illustrates the whole approach. Note that one key advantage of this compared to other TS foundation architectures is that the context length is flexible by design, i.e. the CNN and the attention mechanism may take as input a context signal of arbitrary length.
@@ 3.2 Model training
For training, a collection of just $34$ different chaotic and cyclic DS is used [29], see Appx. A.2 for full specification. From each, multivariate time series $\bm{X}\in\mathbb{R}^{N\times T}$ are simulated ($\approx 1.6$ million in total) and then standardized (dimension-wise), of which the first $T_{C}<T$ column entries are defined as the context signal $\bm{C}=\bm{X}_{1:T_{C}}$. The experts are initialized at $t_{0}=T_{C}-\Delta t+1$ and forward-iterated until time $T$, such that they overlap for a period $\Delta t$ with the context. This is to ensure that the model learns to optimally utilize the context for generalization. A MSE loss is then computed across ground truth, $\bm{X}_{T_{C}-\Delta t+1:T}$, and respective model-generated time series $\hat{\bm{X}}$. The model is trained by sparse teacher forcing (STF; [6, 56]), a control-theoretic technique specifically designed for DSR. STF replaces part of the forward-iterated latent states $\bm{z}_{t}$ by data-inferred states $\hat{\bm{z}}_{t}$ at optimally chosen intervals $\tau$, to avoid exploding gradients even for chaotic systems while enabling the DSR model “to explore the future” (see [56] for detailed theoretical motivation). As shown in Fig. 24, STF with optimal $\tau$ is essential for achieving good zero-shot DSR results. Note that STF is only used for training and turned off at test time.
@@ 3.3 DSR evaluation & time series forecasting
We used two established measures to assess DSR quality [46, 6, 38, 89, 93, 66, 56, 61, 30]: First, to quantify agreement in state space (attractor) geometry, we employed a Kullback-Leibler divergence defined across space, $D_{stsp}$ (see Appx. A.1 for details). For assessing agreement in long-term temporal properties, we used the Hellinger distance $D_{H}$ defined across power spectra of the true and model-generated trajectories. In addition, a short-term, $n$-step-ahead mean absolute prediction error was evaluated. As has been emphasized repeatedly in the statistical and DSR literature [89, 46, 6, 56, 64, 30], prediction errors defined on time series are often only sensible on short time scales, because of the well known exponentially fast divergence of trajectories in chaotic DS (see Fig. 22).
@@ 4.1 Zero-shot DSR
Figure 3: a) DynaMix zero-shot DSR (red) compared to ground truth (blue) using a $500$-step context (gray) for the Sprott M system. b) Zero-shot forecasts (red) for the Selkov DS (true vector field in blue) from different initial conditions outside the context range (gray). c) DSR quality as a function of context length. d) DSR quality as a function of the temporal resolution $\Delta t$ of the context signal.
@@
Figs. 1 (top) & 3a present example zero-shot DSRs of new chaotic DS, which were not part of DynaMix’ training domain, from just a provided context signal $\bm{C}$ (see Figs. 10-14 for further examples). Despite the inherent challenge of predicting chaotic systems, where small differences in initial conditions lead to rapidly diverging trajectories (Fig. 22), DynaMix successfully captures the system’s temporal & geometrical behavior, i.e. accurately reconstructs the underlying state space and faithfully recovers the true system’s power spectrum. Fig. 3b highlights further aspects of DynaMix’ reconstructions: Even though not trained on any $2d$ systems, the model successfully recreates the limit cycle from the $2d$ Selkov system [73]. Notably, the model even generalizes to new initial conditions outside the scope of the context data, i.e. correctly infers properties of the state space beyond the short context trajectory. As further shown in Fig. 3c, the context length necessary to achieve good DSR quickly converges, with as few as $500$ time steps often sufficient. Fig. 3d illustrates that our approach also works across a surprisingly wide range (about an order of magnitude) of temporal resolutions at which the underlying DS was sampled, confirming its robustness w.r.t. sampling frequency.
@@
We next compared the zero-shot, out-of-domain DSR performance of DynaMix to that of a number of recent TS foundation models, including various versions of Chronos [2], Mamba4Cast [5], and Tiny Time Mixers [26], using the long-term statistics $D_{stsp}$ and $D_{H}$ for evaluation (see sect. A.1). As seen in Fig. 4a-b, DynaMix clearly and significantly outperforms all of these. In terms of short-term ($10$-step ahead) forecasts, for which TS foundation models (unlike DynaMix) are optimized, DynaMix performs on par with the (strongest in our batch) Chronos-t5 models (t-tests, all $p>0.37$), as shown in Fig. 4c. Table LABEL:tab:performance_DSR assembles further results on performance comparisons, while Figs. 15-20 provide specific examples. Moreover, a significant limitation of particularly the large Chronos models with millions of parameters is their extreme inefficiency in both inference time and computational costs, as compared in Fig. 5. In fact, their inference time can even exceed that of a custom-trained model. In contrast, DynaMix provides a highly efficient alternative, offering superior DSR performance while using much less parameters and computation time.
Figure 4: Zero-shot DSR performance across all 54 test set DS for DynaMix and various TS foundation models for context length $T_{C}=2000$ (see Fig. 21 for results with $T_{C}=512$). Median±$\pm$±MAD of a) $D_{stsp}$ (geometrical disagreement), b) $D_{H}$ (temporal disagreement), and c) short-term prediction error.
@@
Figure 5: Inference times for zero-shot generation of Lorenz-63 DS with $10000$ forecasting steps. All models were provided the exact same context and run on the same hardware. Note the log-scale on the $x$-axis. For comparison, also the time needed for training and inference of a custom-trained AL-RNN is shown in red.
@@ 4.2 Reasons for the failure of TS foundation models on DSR
Why do TS foundation models perform so poorly on DSR problems? Essentially, they face three key issues: First, most TS foundation models cannot efficiently deal with multivariate time series, but treat dimensions independently. However, in a nonlinear DS all variables are usually coupled, and their joint evolution is actually defining for the system’s (long-term) dynamics and attractor states. Thus, ignoring this, as illustrated in Fig. 6a, results in poor long-term accuracy (see also Figs. 15-20). This highlights a fundamental limitation: Univariate approaches are inherently insufficient for reconstructing multivariate, interacting DS. Second, TS foundation models are trained for short-term prediction, but not for DSR, providing another reason why they fail to reconstruct the correct long-term behavior. As illustrated in Figs. 1 & 6b, in the longer run they often converge to fixed points or simple cycles where the true dynamics is chaotic, not reflecting the context anymore. Third, context parroting, a phenomenon originally identified in LLMs [4], has also been observed in TS foundation models [93], where they tend to repetitively reproduce the exact input pattern rather than dynamically adapting to the system’s true evolution. Thus, TS foundation models simply produce oscillations, where the true dynamics is chaotic and therefore by definition irregular and non-repeating (Fig. 6c). This is illustrated in more detail in Fig. 6d for Chronos-t5-base, the best of the TS foundation models in our batch. It is particularly evident in the power spectrum (Fig. 6d right), where Chronos’ sharp, narrow peaks wrongly suggest periodicity, while the true spectrum is usually rather broad and smeared out for chaotic systems, as the one exhibited by DynaMix. We further quantified this by calculating from the forecast time series the maximum Lyapunov exponent [56], a measure for the exponential divergence rate of trajectories (see Appx. A.1). For Chronos, the value $\lambda_{Chronos}\approx 0.06$ close to zero confirms the nearly periodic behavior, while DynaMix produces exponentially diverging trajectories with $\lambda_{DynaMix}\approx 1.48$ close to the ground truth value of $\lambda_{Lorenz63}\approx 1.37$. Hence, unlike DynaMix, Chronos does not reconstruct the underlying dynamics but just repeats the context. On the other hand, if the underlying dynamics truly is cyclic, DynaMix settles into a cyclic pattern as well, exhibiting no intrinsic bias, as illustrated in Fig. 6e.
@@
Figure 6: Common problems in TS foundation models (red: forecast, gray: context, blue: ground truth). a) Dimensions are dynamically decoupled. b) Long-term forecasts often converge to simple fixed points or cycles unrelated to the true dynamics. c) Context parroting covers up the true dynamics. d) Example forecasts for the chaotic Lorenz-63 system with DynaMix and Chronos-t5-base. Chronos’ context parroting results in cyclic repetition of a fixed pattern, thus inherently missing the true aperiodic, chaotic dynamics. This is evident in the close-to-0 max. Lyapunov exponent (center) and peaked power spectrum (right). e) For comparison, if the underlying behavior is truly cyclic with prominent peaks in the power spectrum, both Chronos and DynaMix are able to capture it.

### v2

@@ Abstract
Complex, temporally evolving phenomena, from climate to brain activity, are governed by dynamical systems (DS). DS reconstruction (DSR) seeks to infer generative surrogate models of these from observed data, reproducing their long-term behavior. Existing DSR approaches require purpose-training for any new system observed, lacking the zero-shot and in-context inference capabilities known from LLMs. Here we introduce DynaMix, a novel multivariate ALRNN-based mixture-of-experts architecture pre-trained for DSR, the first DSR model able to generalize zero-shot to out-of-domain DS. Just from a provided context signal, without any re-training, DynaMix faithfully forecasts the long-term evolution of novel DS where existing time series (TS) foundation models, like Chronos, fail – at a fraction of the number of parameters ($0.1\%$) and orders of magnitude faster inference times. DynaMix outperforms TS foundation models in terms of long-term statistics, and often also short-term forecasts, even on real-world time series, like traffic or weather data, typically used for training and evaluating TS models, but not at all part of DynaMix’ training corpus. We illustrate some of the failure modes of TS models for DSR problems, and conclude that models built on DS principles may bear a huge potential also for advancing the TS prediction field.
@@ 1 Introduction
  • Accurate zero-shot DSR: DynaMix achieves strong generalization across diverse DS, eliminating the need for fine-tuning while maintaining accuracy in reproducing attractor geometry and long-term statistics.
  • Multivariate information transfer: Due to its multivariate architecture, the model efficiently captures dependencies among multiple system dimensions, enabling accurate reconstruction of their coupled dynamics. Further, it is neither bound to a specific dimensionality nor to a specific context length, but can flexibly adapt to other dimensions through specific embeddings.
  • Computational and parameter efficiency: The model achieves high performance with a very lightweight architecture ($\approx 10$k parameters) and small training corpus ($34$ systems), enabling orders of magnitude faster inference than other foundation models.
@@ 2 Related work
Large language models (LLMs) exhibit an impressive ability to infer patterns from prompts (context) and generalize to novel situations without retraining [10, 27, 21, 18], although the underlying reasons for this are still a matter of debate [51, 90, 96]. This inspired the development of general-purpose time series (TS) foundation models which could accurately forecast time series from a short segment provided ‘in-context’ [2, 26], without the need of task-specific fine-tuning [20]. One idea is to simply use pretrained LLMs directly as TS forecasters [34, 71, 79, 93, 83], but the inherent differences between textual and many temporal data, such as continuity, present significant challenges [82]. To overcome these, transformer-based architectures, similar in design to LLMs, such as Chronos [2] or TimesFM [20], were specifically pretrained on a large corpus of time series data. Promising zero-shot forecasting capabilities have also been achieved with alternative architectural designs, such as Tiny Time Mixers [26] or state-space models like Mamba [5].
@@ 3.1 Model architecture
To enable zero-shot reconstruction of novel DS, we develop a specific mixture-of-experts (MoE) architecture that can be pretrained across many diverse DS (Fig. 10), with different experts possibly specializing in different dynamical regimes. As a SOTA DSR base model for the experts, we leverage a recent parameter-friendly RNN, which allows for highly efficient DSR training and is designed to yield topologically parsimonious and interpretable representations of DS it is being trained on, the Almost-Linear RNN (AL-RNN; [7]):
@@
The selection of AL-RNN experts is achieved through a gating network (Fig. 2), which receives as inputs the (generally multivariate) context time series $\bm{C}=\{\bm{c}_{t}\}\in\mathbb{R}^{N\times T_{C}}$ as well as the current latent state $\bm{z}_{t}$. Both are passed into a state attention mechanism defined as
$\bm{w}^{att}_{t}=\sigma\left(\frac{\left|\bm{C}-\left(\bm{Dz}_{t}+\bm{\epsilon}\right)\bm{1}_{T_{C}}^{\top}\right|^{\top}\bm{1}_{N}}{\tau_{\text{att}}}\right)\in\mathbb{R}^{T}\;,$ (3)
where $\bm{D}\in\mathbb{R}^{N\times M}$ is a learnable matrix which maps the latent state into observation space, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{\Sigma})$ exploration noise with a learnable covariance $\bm{\Sigma}$, $\tau_{\text{att}}$ a learnable temperature parameter, $\bm{1}_{\{T_{C},N\}}$ column vectors of ones of length $T_{C},N$, respectively, and $\sigma(\cdot)$ the softmax returning normalized weights. Hence, this mechanism computes attention weights based on some distance between projected latent states and actual context observations. At the same time, the context signal $\bm{C}$ is processed by a CNN, yielding temporal features $\tilde{\bm{C}}\in\mathbb{R}^{N\times T_{C}}$. These features are then weighted with the attention weights $\bm{w}^{att}_{t}$, and together with the current state $\bm{z}_{t}$ further processed by a multi-layer perceptron (MLP) followed by a softmax, yielding a set of expert weights at any time $t$:
$\bm{w}^{exp}_{t}=\sigma\left(\frac{\text{MLP}(\tilde{\bm{C}}{\bm{w}^{att}_{t}},\bm{z}_{t})}{\tau_{\text{exp}}}\right)\in\mathbb{R}^{J}\;,$ (4)
where $\tau_{\text{exp}}$ is another learnable temperature parameter. The forward-iterated latent states $\bm{z}_{t+1}^{j}$ of the individual experts $j\in\{1,...,J\}$ are then weighted by $\bm{w}^{exp}_{t}$ to yield the next time step prediction $\bm{z}_{t+1}=\sum_{j=1}^{J}w^{exp}_{j,t}\cdot\bm{z}_{t+1}^{j}$. Fig. 2 illustrates the whole approach. Note that one key advantage of this compared to other TS foundation architectures is that the context length is flexible by design, i.e. the CNN and the attention mechanism may take as input a context signal of arbitrary length.
@@ 3.2 Model training
For training, a collection of just $34$ different chaotic and cyclic DS is used [29], see Appx. A.2 for full specification and Fig. 10 for illustration. From each, multivariate time series $\bm{X}\in\mathbb{R}^{N\times T}$ are simulated ($\approx 6\times 10^{5}$ in total) and then standardized dimension-wise (likewise for the test set), of which the first $T_{C}<T$ column entries are defined as the context signal $\bm{C}=\bm{X}_{1:T_{C}}$. The experts are initialized at $t_{0}=T_{C}-\Delta t+1$ and forward-iterated until time $T$, such that they overlap for a period $\Delta t$ with the context. This is to ensure that the model learns to optimally utilize the context for generalization. A MSE loss is then computed across ground truth, $\bm{X}_{T_{C}-\Delta t+1:T}$, and respective model-generated time series $\hat{\bm{X}}$. The model is trained by sparse teacher forcing (STF; [6, 58]), a control-theoretic technique specifically designed for DSR. STF replaces part of the forward-iterated latent states $\bm{z}_{t}$ by data-inferred states $\hat{\bm{z}}_{t}$ at optimally chosen intervals $\tau$, to avoid exploding gradients even for chaotic systems while enabling the DSR model “to explore the future” (see [58] for detailed theoretical motivation). As shown in Fig. 26, STF with optimal $\tau$ is essential for achieving good zero-shot DSR results. Note that STF is only used for training and turned off at test time.
@@ 3.3 DSR evaluation & time series forecasting
We used two established measures to assess DSR quality [48, 6, 39, 94, 99, 70, 58, 64, 30]: First, to quantify agreement in state space (attractor) geometry, we employed a Kullback-Leibler divergence defined across space, $D_{stsp}$ (see Appx. A.1 for details). For assessing agreement in long-term temporal properties, we used the Hellinger distance $D_{H}$ defined across power spectra of the true and model-generated trajectories. In addition, a short-term, $n$-step-ahead mean absolute prediction error (MAE) was evaluated. As has been repeatedly emphasized in the statistical and DSR literature [94, 48, 6, 58, 68, 30], prediction errors defined on time series are often only sensible on short time scales, because of the well known exponentially fast divergence of trajectories in chaotic DS (see Fig. 24).
@@ 4.1 Zero-shot DSR
Figure 3: a) DynaMix zero-shot DSR (red) compared to ground truth (lightgray) using a $500$-step context (darkgray) for the Sprott M system. b) Zero-shot forecasts for the Selkov DS (true vector field in lightgray) from different initial conditions (red) outside the context range (darkgray). c) DSR quality as a function of context length for Lorenz-63. d) DSR quality as a function of the temporal resolution $\Delta t$ of the context signal. Error bands = STD
@@
Figs. 1 (top) & 3a present example zero-shot DSRs of new chaotic DS, which were not part of DynaMix’ training domain, from just a provided context signal $\bm{C}$ (see Figs. 11-15 for further examples). Despite the inherent challenge of predicting chaotic systems, where small differences in initial conditions lead to rapidly diverging trajectories (Fig. 24), DynaMix successfully captures the system’s temporal & geometrical behavior, i.e. accurately reconstructs the underlying state space and faithfully recovers the true system’s power spectrum. Fig. 3b highlights further aspects of DynaMix’ reconstructions: Even though not trained on any $2d$ systems, the model successfully recreates the limit cycle from the $2d$ Selkov system [77]. Notably, the model even generalizes to new initial conditions outside the scope of the context data, i.e. correctly infers properties of the state space beyond the short context trajectory (see Fig. 22 for further examples). Likewise, although trained only on $3d$ DS, DynaMix also successfully generalizes to higher-dimensional DS, as illustrated for the $6d$ Lorenz-96 system [54] in Fig. 34 (see Appx. A.6 for details on the setup and comparisons to other models). As further shown in Fig. 3c, the context length necessary to achieve good DSR quickly converges, with as few as $500$ time steps often sufficient. Fig. 3d illustrates that our approach also works across a surprisingly wide range (about an order of magnitude) of temporal resolutions at which the underlying DS was sampled, confirming its robustness w.r.t. sampling frequency.
@@
We next compared the zero-shot, out-of-domain DSR performance of DynaMix to that of a number of recent TS foundation models, including various versions of Chronos [2], Mamba4Cast [5], TimesFM [20], and Tiny Time Mixers [26], using the long-term statistics $D_{stsp}$ and $D_{H}$ for evaluation (see Sect. A.1). As seen in Fig. 4a-b, DynaMix clearly and significantly outperforms all of these. In terms of short-term ($10$-step ahead) forecasts, for which TS foundation models (unlike DynaMix) are optimized, DynaMix performs on par with the (strongest in our batch) Chronos-t5 models, as shown in Fig. 4c. Table LABEL:tab:performance_DSR assembles further results on performance comparisons, while Figs. 16-21 provide specific examples. Moreover, a significant limitation of particularly the large Chronos models with millions of parameters is their extreme inefficiency in both inference time and computational costs, as compared in Fig. 5. In fact, their inference time can even exceed that of a custom-trained model. In contrast, DynaMix provides a highly efficient alternative, offering superior DSR performance while using orders of magnitude less parameters ($\approx 10$k in total) and computation time.
[…]
@@
Figure 5: Number of model parameters vs. inference times for zero-shot generation of Lorenz-63 DS with $10000$ forecasting steps. All models were provided the exact same context and run on the same hardware. Note the log-scale on the xx/$y$-axis. For comparison, also the time needed for training and inference of a custom-trained AL-RNN is shown (turquoise).
To put DynaMix’ performance further in context, we also compared it to a couple of custom-trained DSR models, i.e. SOTA DSR models explicitly trained on the context data, incl. the plain AL-RNN trained by STF [7], Neural ODEs [37, 17], and reservoir computers which are commonly used for DSR [70] (see Appx. A.7 for details). Despite the custom trained models’ unduly advantage, performing – unlike DynaMix – only in-domain generalization [37], DynaMix keeps up with, and in some cases even outperforms, them (Appx. Tables 68).
@@ 4.2 Reasons for the failure of TS foundation models on DSR
Why do TS foundation models perform so poorly on DSR problems? Essentially, they face three key issues: First, most TS foundation models cannot efficiently deal with multivariate time series, but treat dimensions independently. However, in a nonlinear DS all variables are usually coupled, and their joint evolution is actually defining for the system’s (long-term) dynamics and attractor states. Thus, ignoring this, as illustrated in Fig. 6a, results in poor long-term accuracy (see also Figs. 16-21). This highlights a fundamental limitation: Univariate approaches are inherently insufficient for reconstructing multivariate, interacting DS. Second, TS foundation models are trained for short-term prediction, but not for DSR, providing another reason why they fail to reconstruct the correct long-term behavior. As illustrated in Figs. 1 & 6b, in the longer run they often converge to fixed points or simple cycles where the true dynamics is chaotic, not reflecting the context anymore. Third, context parroting, a phenomenon originally identified in LLMs [4], has also been observed in TS foundation models [99], where they tend to repetitively reproduce the exact input pattern rather than dynamically adapting to the system’s true evolution. Thus, TS foundation models simply produce oscillations, where the true dynamics is chaotic and therefore by definition irregular and non-repeating (Fig. 6c).
@@
Figure 6: Common problems of TS foundation models (red: forecast, darkgray: context, lightgray: ground truth). a) Dimensions are dynamically decoupled. b) Long-term forecasts often converge to simple fixed points or cycles unrelated to the true dynamics. c) Context parroting covers up the true dynamics. d) Example forecasts for the chaotic Lorenz-63 system with DynaMix and Chronos-t5-base. Chronos’ context parroting results in cyclic repetition of a fixed pattern, thus inherently missing the true aperiodic, chaotic dynamics. This is evident in the close-to-0 max. Lyapunov exponent (center) and peaked power spectrum (right). e) For comparison, if the underlying behavior is truly cyclic with prominent peaks in the power spectrum, both Chronos and DynaMix are able to capture it.
This is illustrated in more detail in Fig. 6d for Chronos-t5-base, the best of the TS foundation models in our batch. It is particularly evident in the power spectrum (Fig. 6d right), where Chronos’ sharp, narrow peaks wrongly suggest periodicity, while the true spectrum is usually rather broad and smeared out for chaotic systems, as the one exhibited by DynaMix. We further quantified this by calculating from the forecast time series the maximum Lyapunov exponent [58], a measure for the exponential divergence rate of trajectories (see Appx. A.1). For Chronos, the value $\lambda_{Chronos}\approx 0.02$ close to zero confirms the nearly periodic behavior, while DynaMix produces exponentially diverging trajectories with $\lambda_{DynaMix}\approx 0.96$ close to the ground truth value of $\lambda_{Lorenz63}\approx 0.87$. Hence, unlike DynaMix, Chronos does not reconstruct the underlying dynamics but just repeats the context. On the other hand, if the underlying dynamics truly is cyclic, DynaMix settles into a cyclic pattern as well, exhibiting no intrinsic bias, as illustrated in Fig. 6e.